OpenStudy (anonymous):

If the perimeter of a rectangular property is 46 meters, and the area of the property is 76 meters, what is the length of each of the shorter sides?

OpenStudy (anonymous):

4

OpenStudy (anonymous):

\[2x+2y=46\] x*y=76 y=76/x x^2-23x+76=0

OpenStudy (anonymous):

Let w and L be the width and length of the rectangular property. Solve the following for w and L: {2 L + 2 w = 46, L w = 76} {L = 19. w = 4} The shorter sides each have a length of 4 meters.